What would you pay for a $170,000 debenture bond that matures in 15 years and pays $8,500 a year in interest if you wanted to earn a yield of:

3%?

4%?

5%?
Answer #1

1) Solution: 210,588.58

Working:

170,000 (P/F, 3% , 15) + 8,500 (P/A, 3%, 15)

= 170,000 * 0.641862 + 8,500 * 11.937935

= 210,588.98

2) Solution: 188,901.26

Working:

= 170,000 (P/F, 4% , 15) + 8,500 (P/A, 4%, 15)

= 94,506.29 + 94,394.96546

= 188,901.26

3) Solution: 170,000

Working:

= 170,000 (P/F, 5% , 15) + 8,500 (P/A, 5%, 15)

= 88,277.09 + 81,772.90668

= 170,000